# From Source to Target: Leveraging Transfer Learning for Predictive Process Monitoring in Organizations

### v1

@@ 6.1.1 Use case I: Intra-organizational transfer
Intra-organizational TL allows an organization to transfer knowledge across its business processes. To evaluate our TL-based PPM technique in the intra-organizational use case, we applied it to process data handling different products. To create this scenario, we used two sub-logs from one organization, specifically sub-logs of the bpic2014 event log [^14] https://data.4tu.nl/articles/_/12692378/1 (Accessed 23 July 2025). The data were provided by an information and communications technologies (ICT) company, Rabobank Group ICT, from the Netherlands. It concerns their ITSM processes, specifically the incident management process. The data were derived from the company’s ITIL service management tool called the HP Service Manager and refer to incidents that are recorded once a reported problem cannot be resolved over the phone. Each incident is related to a service component, that is, one particular product in the bill of materials. The first sub-log WBS72&223, the source log, only includes traces related to the service components WBS000072 and WBS000223. The second sub-log WBS263, our target log, only includes traces related to the service component WBS000263. In the pre-processing, traces longer than 50 events were removed.
@@
Table 3: Activities and their frequency for both event logs for the intra-organizational use case.
@@
For our evaluation, we predict whether the process instance will be finished in time or not. The label was constructed by taking the 70th quantile of all traces’ duration as the criterion for in-time for each sub-log independently. The value of the criterion for the source log WBS72&223 was 4.94 days, and for the target log WBS263, 4.15 days. Traces with a longer duration were labeled as not in-time.
@@ 6.1.2 Use case II: Inter-organizational transfer
Table 5: Activities and their frequency for both event logs for the inter-organizational use case.
@@
Just like the intra-organizational use case, we predict whether the process instance is in time or not. The label was constructed by taking the 70th quantile of all traces’ duration as the criterion for in-time for each log independently. The value of the criterion for the source log bpic2014 was 2.83 days, and for the target log helpdesk, 44.95 days. Traces with a longer duration were labeled as not in-time.
@@ 6.2 Setup
For our first three evaluations on the prediction performance of our technique, we sorted the process instances of the event log of the source context and the target context in temporal order according to the timestamps of their first event. We then split each event log into a 64% training, a 16% validation, and a 20% test set, trained the models, and computed $\text{AUC}_{\text{ROC}}$ and Matthews correlation coefficient (MCC), as well as the weighted precision, recall, and F1-score based on the prediction results on test sets. Among these ML metrics, we choose the $\text{AUC}_{\text{ROC}}$ as the primary metric because it is independent of the threshold, which, therefore, allows for a more holistic evaluation of the models’ prediction performance. We repeated each experiment five times with different seeds and calculated the average values and the standard deviation across the five runs. In addition, for each approach tested in the first three evaluations, we used two additional baselines and compared their prediction performance with the model of the respective approach trained on the event data of the source context and transferred to the target context without fine-tuning. The first baseline is a model trained and tested on the event data of the source context, which serves as a model with a reference prediction performance in the source context. The second baseline is a model trained and tested on the event data of the target context, which serves as a model with a reference prediction performance in the target context.
@@
For the first evaluation, in which we compared the prediction performance of our TL-based technique to that of the traditional PPM approaches, we selected five baseline approaches. The first baseline employs the same LSTM model as in our proposed technique. To encode the activities for this and the other models, one-hot encoding is applied. This is a common technique for encoding categorical attributes of events in PPM (e.g., Rama-Maneiro et al. 2023). However, since we do not know the target context’s activities when training the base model in the source context, only the source context’s activities are used to fit the one-hot encoder. The encoder is applied to transform the target context’s activities. Consequently, if an activity exists in both contexts, it can be encoded in the target context’s event data; otherwise, this is not possible. In addition, the duration since start time feature is min-max scaled using the training event data from the source context. The second baseline employs a logistic regression model with the same encoding strategy as the first approach. The last three baselines use either an extreme gradient boosting (XGBoost), a random forest, or a decision tree model. The activities are one-hot encoded again, while the duration since start time feature is not scaled or normalized. This is because the underlying algorithms for building these models are tree-based and split data based on feature thresholds rather than the absolute magnitude or distribution of features. The models employed in all baselines are commonly used in outcome-oriented PPM (Teinemaa et al. 2019; Maggi et al. 2014; Wang et al. 2019).
@@
For the second evaluation, in which we investigated the effect of using different pre-trained models for activity encoding on prediction performance, three types of embedding models were employed. The first group includes three static word embedding models. The first two models, glove-wiki-gigaword-100 and glove-wiki-gigaword-300, are GloVe models [^17] https://radimrehurek.com/gensim/models/word2vec.html#other-embeddings (Accessed 23 July 2025) that output feature vectors of length 100 and 300 for the activities, respectively. The third model, word2vec-google-news-300, is a pre-trained word2vec model [^18] https://radimrehurek.com/gensim/models/word2vec.html (Accessed 23 July 2025) that encodes activities as vectors of length 300. Although these encoding models have the limitation that all input words must be included in the vocabulary of the pre-trained model, we did not replace any domain-specific acronyms included in the activity names to ensure a fair comparison. For example, we did not replace the acronym “CI” with “configuration item” in the activity name “Caused By CI” of the bpic2014 event log. The second group includes two contextual word embedding models. The first model is bert-base-cased [^19] https://huggingface.co/google-bert/bert-base-cased (Accessed 23 July 2025), a pre-trained BERT model (Devlin et al. 2019). This BERT model has the same architecture as the original BERT model, which is entirely based on the encoder stack of the transformer model (Vaswani et al. 2017). Pre-training was performed on a large corpus using masked language modeling (prediction of randomly masked words in sentences based on the surrounding words) and next sentence prediction objectives. The second model is bert-base-uncased [^20] https://huggingface.co/google-bert/bert-base-uncased (Accessed 23 July 2025), another pre-trained BERT model, which differs from the bert-base-cased model by lower-casing all input text before tokenization. Both models output contextualized word embeddings of length 748 for each token. As activities typically consist of several tokens, a mean pooling operation is performed on top of the contextual word embeddings to produce one embedding vector per activity. The third group includes two contextual sentence embedding models. The first model is all-MiniLM-L12-v2212121https://huggingface.co/sentence-transformers/all-MiniLM-L12-v2 (Accessed 23 July 2025), a fine-tuned version of the pre-trained Mini-LM model (Wang et al. 2020) MiniLM-L12-H384-uncased [^22] https://huggingface.co/sentence-transformers/all-MiniLM-L12-v2 (Accessed 23 July 2025). Fine-tuning is performed using a contrastive learning objective (predicting for a sentence the correct matching sentence from a group of randomly selected sentences), which allows for a better capture of the semantic meaning of entire sentences. The second model is all-mpnet-base-v2232323https://huggingface.co/sentence-transformers/all-mpnet-base-v2 (Accessed 23 July 2025), a version of the pre-trained mpnet model (Song et al. 2020) for which fine-tuning was also performed using a contrastive learning objective. The all-MiniLM-L12-v2 and all-mpnet-base-v2 models output contextualized word embeddings of lengths 384 and 768, respectively, and, as with the contextual word embedding models, a mean pool operation was performed to get one embedding vector per activity.
@@
For the fourth evaluation, in which we assessed the effect of the proposed relative cross-domain approach for timestamp encoding on the prediction performance, we employed two different baseline approaches. In the first approach, the time $h$ (in the form of the number of hours), the day of the week $d$, and the month $m$ are extracted from the timestamp and coded as separate features. These variables have a cyclical structure: a Sunday is followed by a Monday, December is followed by January, and the 23rd hour is followed by the 0-th hour. To take this periodic nature into account, the coding is carried out using a sinus function that can depict cyclical relationships. For this purpose, the months are numbered from 0 (January) to 11 (December), the days of the week from 0 (Monday) to 6 (Sunday), and the hours from 0 to 23, and transformed as follows:
@@ 7.1 Use case I: Intra-organizational transfer
For the intra-organizational use case, we evaluated the prediction performance of our technique by comparing it with traditional PPM approaches. In addition, we investigated the feature-based transfer of our technique in terms of different activity embeddings, including the quality of embeddings, and different timestamp encodings. Lastly, the transfer relevance was examined.
@@ 7.1.2 Prediction performance for different activity embedding models
To find an effective activity embedding model for our proposed TL-based technique for the intra-organizational use case, we explored seven activity embedding models. Table 7 shows the results of these seven embedding models, which can be divided into three categories: i) static word, ii) contextual word, and iii) contextual sentence embedding models. Embedding the activity feature with the contextual sentence embedding model all-mpnet-base-v2 achieves the best results in all metrics, that is, precision, recall, F1-score, MCC, and $\text{AUC}_{\text{ROC}}$.
@@ 7.1.3 Activity embedding quality
To investigate the feature-based TL part of our technique in more detail, we compared the difference in activity embeddings for the all-mpnet-base-v2 embedding model [^25] The figures showing the Euclidean distance between activity embedding vectors created with the other embedding models of our setting can be found in the online repository. (see Figure 5) between $L_{S}$ and $L_{T}$ by calculating the Euclidean distance between the feature vectors.
The source and target logs have 25 activities in common, which can be seen in Figure 5 as the black diagonal line formed by the 0.00 distances between the activity embeddings. In both event logs, $L_{S}$ and $L_{T}$, there are two groups of activities (indicated by the deep-shaded squares in Figure 5), which contain activities that are recorded at a more granular level. The first group deals with communication, either Communication with customer or Communication with vendor. The second group concerns the quality indicator, either referring to Quality Indicator itself, or providing more details, as in Quality Indicator Fixed or Quality Indicator Set. For both groups, the embeddings of the more granular activities are relatively similar, displaying a smaller distance in embedding vectors.
@@ 7.1.4 Prediction performance for different timestamp encoding approaches
Next to investigating strategies to embed activities of our TL-based PPM technique, we investigated five additional encoding strategies for the time-based feature duration since start time and compared them to our technique’s relative cross-domain encoding. These five encoding strategies fall into two categories: i) time-based encoding and ii) autoencoder-based encoding. The results in Table 8 indicate a clear trend towards the relative cross-domain encoding as it achieves the highest precision, recall, F1-score, MCC, and $\text{AUC}_{\text{ROC}}$ out of the six encoding strategies. The second-best encoding in all metrics is the autoencoder-based encoding.
@@ 7.1.5 Transfer relevance
To provide an overview of how much target data should be available to train a model outperforming our transferred model (not fine-tuned), Figure 6 shows the $\text{AUC}_{\text{ROC}}$ for training a model solely on event data available in the target context with the first 1%, 2%, 5%, 10%, 20%, 50%, and 100% of cases using the all-mpnet-base-v2 (see Section 7.1.2) embedding model for activity encoding and the relative cross-domain approach (see Section 7.1.4) for timestamp encoding.
@@ 7.2 Use case II: Inter-organizational transfer
The traditional PPM technique that employs the XGBoost model attains a higher recall, while the approach with the random forest model yields a superior MCC. Comparing the results of the transferred models with the baseline approach Train and test on $L_{T}$, transferred models come close to the prediction performance of the baseline approach Train and test on $L_{T}$ in most metrics.
@@ 7.2.3 Activity embedding quality
Figure 7 shows the Euclidean distance between the activity embedding vectors from $L_{S}$ and $L_{T}$, created via the glove-wiki-gigaword-100 embedding model. [^26] The figures showing the Euclidean distance between activity embedding vectors created with the other embedding models of our setting can be found in the online repository.
The source and target logs share only two common activities, Closed and RESOLVED, which are represented by an activity embedding distance of 0.00, as shown in Figure 5.
The other twelve activities in $L_{T}$ cannot be found in the source log $L_{S}$. Interestingly, there is no clear match based on the smallest Euclidean distance between the embedding vectors for any of the 12 activities in $L_{T}$ with any of the activities in $L_{S}$. However, based on the embedding distance, these twelve activities can be categorized into two groups. The first group contains seven activities, which have an overall closer match to other activities in $L_{S}$. These are Create SW anomaly, Insert ticket, Require upgrade, Resolve SW anomaly, Resolve ticket, Schedule intervention, and Take in charge ticket. The other five activities in $L_{T}$, Assign seriousness, DUPLICATE, INVALID, VERIFIED, and Wait, are overall further apart from activities in $L_{S}$ based on the Euclidean distance between embedding vectors. However, the activities written in capital letters are outliers in terms of their frequency occurring in $L_{T}$ (see Table 5 in Section 6.1.2).
@@ 7.2.4 Prediction performance for different timestamp encoding approaches
Similar to the intra-organizational use case, we also examined five additional encoding strategies for the time-based feature duration since start time, belonging to the categories time-based encoding and autoencoder-based encoding, and compared them to our technique’s relative cross-domain encoding. The results in Table 8 indicate a clear trend towards the relative cross-domain encoding as it achieves the highest precision, recall, F1-score, MCC, and $\text{AUC}_{\text{ROC}}$ out of the six encoding strategies. The prediction performance of our technique using the time-based and autoencoder-based encoding strategies is comparably lower.
Table 11: Prediction performance of our TL-based PPM technique with the proposed cross-domain encoding of the elapsed time attribute and other encoding approaches for the inter-organizational use case (average and standard deviation over five runs).

### v2

@@ 6.1.1 Use case I: Intra-organizational transfer
Intra-organizational TL allows an organization to transfer knowledge across its business processes. To evaluate our TL-based PPM technique in the intra-organizational use case, we applied it to process data handling different products. In particular, we used the event data of two sub-logs from one organization, specifically sub-logs of the bpic2014 event log [^14] https://data.4tu.nl/articles/_/12692378/1 (Accessed 23 July 2025). The data were provided by an information and communications technologies (ICT) company from the Netherlands, called Rabobank Group ICT. It concerns their ITSM processes, specifically their incident management process. The data were derived from the company’s ITIL service management tool, called the HP Service Manager, and refer to incidents that are recorded when a reported problem cannot be resolved over the phone. Each incident is related to a service component, that is, one particular product in the bill of materials. The first sub-log WBS72&223, the source log, only includes traces related to the service components WBS000072 and WBS000223. The second sub-log WBS263, our target log, only includes traces related to the service component WBS000263. In the pre-processing, traces longer than 50 events were removed.
@@
Table 3: Activities and their frequency of the event logs used for the intra-organizational use case.
@@
For our evaluation, we predict whether running process instances will be finished in time or not. The label was constructed by taking the 70th quantile of all traces’ duration as the criterion for in-time for each sub-log independently. The value of the criterion for the source log WBS72&223 was 4.94 days, and for the target log WBS263, 4.15 days. Traces with a longer duration were labeled as not in-time.
@@ 6.1.2 Use case II: Inter-organizational transfer
Table 5: Activities and their frequency of the event logs used for the inter-organizational use case.
@@
Just like the intra-organizational use case, we predict whether running process instances will be finished in time or not. The label was constructed by taking the 70th quantile of all traces’ duration as the criterion for in-time for each log independently. The value of the criterion for the source log bpic2014 was 2.83 days, and for the target log helpdesk, 44.95 days. Traces with a longer duration were labeled as not in-time.
@@ 6.2 Setup
For our first three evaluations on the prediction performance of our technique, we sorted the process instances of the event logs of the source and target contexts in temporal order according to the timestamps of their first event. We then split each event log into a 64% training, a 16% validation, and a 20% test set, trained the models, and computed the $\text{AUC}_{\text{ROC}}$ and the Matthews correlation coefficient (MCC), as well as the weighted precision, recall, and F1-score based on the prediction results on test sets. Among these ML metrics, we choose the $\text{AUC}_{\text{ROC}}$ as the primary metric because it is independent of the threshold, which, therefore, allows for a more holistic evaluation of the models’ prediction performance. We repeated each experiment five times with different seeds and calculated the average values and the standard deviation across the five runs. In addition, for each approach tested in the first three evaluations, we used two additional baselines and compared their prediction performance with the model of the respective approach trained on the event data of the source context and transferred to the target context without fine-tuning. The first baseline is a model trained and tested on the event data of the source context, which serves as a model with a reference prediction performance in the source context. The second baseline is a model trained and tested on the event data of the target context, which serves as a model with a reference prediction performance in the target context.
@@
For the first evaluation, in which we compared the prediction performance of our TL-based technique to that of the traditional PPM approaches, we selected five baseline approaches. The first baseline employs the same LSTM model as in our proposed technique. To encode the activities for this and the other models, one-hot encoding is applied. It is a common technique for encoding categorical attributes of events in PPM (e.g., Rama-Maneiro et al. 2023). However, since we do not know the target context’s activities when training the base model in the source context, only the source context’s activities are used to fit the one-hot encoder. The encoder is then applied to transform the target context’s activities. Consequently, if an activity exists in both contexts, it can be encoded in the target context’s event data; otherwise, this is not possible. In addition, for the duration since start time feature, a min-max scaler is only fitted on the training event data from the source context and applied to the test event data from the source and target contexts to avoid data leakage. The second baseline employs a logistic regression model with the same encoding strategy as the first approach. The last three baselines use either an extreme gradient boosting (XGBoost), a random forest, or a decision tree model. The activities are one-hot encoded again, while the duration since start time feature is not scaled or normalized. This is because the underlying algorithms for building these models are tree-based and split data based on feature thresholds rather than the absolute magnitude or distribution of features. The models employed in all baselines are commonly used in outcome-oriented PPM (Teinemaa et al. 2019; Maggi et al. 2014; Wang et al. 2019).
@@
For the second evaluation, in which we investigated the effect of using different pre-trained models for activity encoding on prediction performance, three types of embedding models were employed. The first group includes three static word embedding models. The first two models, glove-wiki-gigaword-100 and glove-wiki-gigaword-300, are pre-trained global vectors for word representation (GloVe) models [^17] https://radimrehurek.com/gensim/models/word2vec.html#other-embeddings (Accessed 23 July 2025) that output feature vectors of length 100 and 300 for the activities, respectively. The third model, word2vec-google-news-300, is a pre-trained word2vec model [^18] https://radimrehurek.com/gensim/models/word2vec.html (Accessed 23 July 2025) that encodes activities as vectors of length 300. Although these encoding models have the limitation that all input words must be included in the vocabulary of the pre-trained model, we did not replace any domain-specific acronyms included in the activity names to ensure a fair comparison. For example, we did not replace the acronym “CI” with “configuration item” in the activity name “Caused By CI” of the bpic2014 event log. The second group includes two contextual word embedding models. The first model is bert-base-cased [^19] https://huggingface.co/google-bert/bert-base-cased (Accessed 23 July 2025), a pre-trained BERT model (Devlin et al. 2019). This BERT model has the same architecture as the original BERT model, which is entirely based on the encoder stack of the transformer model (Vaswani et al. 2017). The pre-training of this model was performed on a large corpus using the masked language modeling (prediction of randomly masked words in sentences based on the surrounding words) and the next sentence prediction objectives. The second model is bert-base-uncased [^20] https://huggingface.co/google-bert/bert-base-uncased (Accessed 23 July 2025), another pre-trained BERT model, which differs from the bert-base-cased model by lower-casing all input text before tokenization. Both models output contextualized word embeddings of length 748 for each token. As activities typically consist of several tokens, a mean pooling operation is performed on top of the contextual word embeddings to produce one embedding vector per activity. The third group includes two contextual sentence embedding models. The first model is all-MiniLM-L12-v2212121https://huggingface.co/sentence-transformers/all-MiniLM-L12-v2 (Accessed 23 July 2025), a fine-tuned version of the pre-trained Mini-LM model (Wang et al. 2020) MiniLM-L12-H384-uncased [^22] https://huggingface.co/sentence-transformers/all-MiniLM-L12-v2 (Accessed 23 July 2025). Fine-tuning is performed using a contrastive learning objective (predicting for a sentence the correct matching sentence from a group of randomly selected sentences), which allows for a better capture of the semantic meaning of entire sentences. The second model is all-mpnet-base-v2232323https://huggingface.co/sentence-transformers/all-mpnet-base-v2 (Accessed 23 July 2025), a version of the pre-trained mpnet model (Song et al. 2020) for which fine-tuning was also performed using a contrastive learning objective. The all-MiniLM-L12-v2 and all-mpnet-base-v2 models output contextualized word embeddings of lengths 384 and 768, respectively, and, as with the contextual word embedding models, a mean pool operation was performed to get one embedding vector per activity.
@@
For the fourth evaluation, in which we assessed the effect of the proposed relative cross-domain approach for timestamp encoding on the prediction performance, we employed two different baseline approaches. In the first approach, the time $h$ (in the form of the number of hours), the day of the week $d$, and the month $m$ are extracted from the timestamp and coded as separate features. These variables have a cyclical structure: Sunday is followed by Monday, December is followed by January, and the 23rd hour is followed by the 0-th hour. To take this periodic nature into account, the coding is carried out using a sinus function that can depict cyclical relationships. For this purpose, the months are numbered from 0 (January) to 11 (December), the days of the week from 0 (Monday) to 6 (Sunday), and the hours from 0 to 23, and transformed as follows:
@@ 7.1 Use case I: Intra-organizational transfer
For the intra-organizational use case, we evaluated the prediction performance of our technique by comparing it with traditional PPM approaches. Furthermore, we investigated the feature-based transfer of our technique in terms of different activity embeddings, including the quality of embeddings, and different timestamp encodings. Lastly, the transfer relevance was examined.
@@ 7.1.2 Prediction performance for different activity embedding models
To find an effective activity embedding model for our proposed TL-based technique for the intra-organizational use case, we explored seven activity embedding models. Table 7 shows the results of these models, which can be grouped into three categories: i) static word, ii) contextual word, and iii) contextual sentence embedding models. Embedding the activity attribute with the contextual sentence embedding model all-mpnet-base-v2 achieves the best results in all metrics, that is, precision, recall, F1-score, MCC, and $\text{AUC}_{\text{ROC}}$.
@@ 7.1.3 Activity embedding quality
To investigate the feature-based TL part of our technique in more detail, we compared the difference in activity embeddings for the all-mpnet-base-v2 embedding model [^25] The figures showing the Euclidean distance between activity embedding vectors created with the other embedding models of our setting can be found in the online repository of this work. between $L_{S}$ and $L_{T}$ by calculating the Euclidean distance between the embedding vectors (see Figure 5).
The source and target logs have 25 activities in common, which can be seen by the black diagonal line formed by the 0.00 distances between the activity embeddings. In addition, in both event logs, $L_{S}$ and $L_{T}$, there are two groups of activities (indicated by the deep-shaded squares), which contain activities that are recorded at a more granular level. The first group deals with communication, either Communication with customer or Communication with vendor. The second group concerns the quality indicator, either referring to Quality Indicator itself, or providing more details, as in Quality Indicator Fixed or Quality Indicator Set. For both groups, the embeddings of the more granular activities are relatively similar, displaying a smaller distance in embedding vectors.
@@ 7.1.4 Prediction performance for different timestamp encoding approaches
Besides investigating strategies to embed activities of our TL-based PPM technique, we examined five additional encoding strategies for the time-based feature duration since start and compared them to our technique’s relative cross-domain encoding. These five encoding strategies fall into two categories: i) time-based encoding and ii) autoencoder-based encoding. The results in Table 8 indicate a clear trend towards the relative cross-domain encoding as it achieves the highest precision, recall, F1-score, MCC, and $\text{AUC}_{\text{ROC}}$ out of the six encoding strategies. The encoding strategy resulting in the second-best prediction performance values in all metrics is the autoencoder-based encoding.
@@ 7.1.5 Transfer relevance
To provide an overview of how much target data should be available to train a model outperforming our transferred model (not fine-tuned), Figure 6 shows the $\text{AUC}_{\text{ROC}}$ for training a model solely on event data available in the target context with the first 1%, 2%, 5%, 10%, 20%, 50%, and 100% of traces using the all-mpnet-base-v2 embedding model (see Section 7.1.2) for activity encoding and the relative cross-domain approach (see Section 7.1.4) for timestamp encoding.
@@ 7.2 Use case II: Inter-organizational transfer
The traditional PPM technique that employs the XGBoost model attains a higher recall, whereas the approach with the random forest model yields a superior MCC. Comparing the results of the transferred models with the baseline approach Train and test on $L_{T}$, the transferred models come close to the prediction performance of the baseline approach Train and test on $L_{T}$ in most metrics.
@@ 7.2.3 Activity embedding quality
Figure 7 shows the Euclidean distance between the embedding vectors of the activities from $L_{S}$ and $L_{T}$, created via the glove-wiki-gigaword-100 embedding model. [^26] The figures showing the Euclidean distance between activity embedding vectors created with the other embedding models of our setting can be found in the online repository of this work.
The source and target logs share only two common activities, Closed and RESOLVED, which are represented by an activity embedding distance of 0.00. The other twelve activities in $L_{T}$ cannot be found in the source log $L_{S}$. Interestingly, there is no clear match based on the smallest Euclidean distance between the embedding vectors for any of the 12 activities in $L_{T}$ with any of the activities in $L_{S}$. However, based on the embedding distance, these twelve activities can be categorized into two groups. The first group contains seven activities, which have an overall closer match to other activities in $L_{S}$. These are Create SW anomaly, Insert ticket, Require upgrade, Resolve SW anomaly, Resolve ticket, Schedule intervention, and Take in charge ticket. The other five activities in $L_{T}$, Assign seriousness, DUPLICATE, INVALID, VERIFIED, and Wait, are overall further apart from activities in $L_{S}$ based on the Euclidean distance between the embedding vectors. However, the activities written in capital letters are outliers in terms of their frequency occurring in $L_{T}$ (see Table 5 in Section 6.1.2).
@@ 7.2.4 Prediction performance for different timestamp encoding approaches
Similar to the intra-organizational use case, we also examined five additional encoding strategies for the time-based feature duration since start, belonging to the categories time-based encoding and autoencoder-based encoding, and compared them to our technique’s relative cross-domain encoding. The results in Table 8 indicate a clear trend towards the relative cross-domain encoding as it achieves the highest precision, recall, F1-score, MCC, and $\text{AUC}_{\text{ROC}}$ out of the six encoding strategies. The prediction performance of our technique using the time-based and autoencoder-based encoding strategies is comparably lower.
Table 11: Prediction performance of our TL-based PPM technique with the proposed cross-domain encoding of the duration since start time feature and other encoding approaches for the inter-organizational use case (average and standard deviation over five runs).